The ticket sales at a movie theater were $2,878. Adult tickets are $11, and senior tickets are $8. The number of senior tickets sold was 25 less than twice the number of adult tickets. Determine the number of adult tickets and number of senior tickets sold.
They sold adult tickets.
They sold senior tickets.

Found 2 solutions by Theo, josgarithmetic:
Answer by Theo(13342)

total sales were 2878.
x = number of adult tickets.
y = number of senior tickets.
adult tickets are 11 each and seniot tickets are 8 each.
equation for that is:
11x + 8y = 2878
number of senior tickets sold was 25 less than twice the number of adult tickets.
equation for that is:
y = 2x - 25
replace y with 2x - 25 in the first equation to get:
11x + 8y = 2878 becomes:
11x + 8 * (2x - 25) = 2878
simplify to get:
11x + 16x - 200 = 2878
combine like terms and add 200 to both sides of the equation to get:
27x = 3078.
solve for x to get:
x = 114.
siince y = 2x - 25, then y = 203.
11x + 8y = 2878 becomes 11*114 + 8*203 = 2878, confirming that the values for x and y are good.
your solution is that they sold 114 adult tickets and 203 senior tickets.
the number of senior tickets is 25 less than twice the number of adult tickets.

Answer by josgarithmetic(39617)

WHICH          PRICE       HOW MANY      MONEY
Adult             11          a          11a
Senior             8        2a-25        8(2a-25)
TOTALS                                   2878

Understand, and see what to do?